OpenStudy (jkbo):

Please help Use the quadratic model y = -4x^2-3x +4 predict y if x equals 5 a.-111 b.-89 c.-51 d.-411

OpenStudy (kainui):

So everywhere you see an x just place a 5 there and calculate it out.

OpenStudy (jkbo):

I got -411 is that correct @Kainui ?

OpenStudy (kainui):

Nope, so try doing each term individually and tell me what they are, then add them together. -4x^2 -3x 4

OpenStudy (jkbo):

the first on is -20 then -15 then 20

OpenStudy (jkbo):

@Kainui

OpenStudy (kainui):

The first term is -4 times 5^2 not 5. Don't forget to square it!

OpenStudy (jkbo):

yeah 4 x 5 is 20

OpenStudy (jkbo):

-20^2 is 400

OpenStudy (jkbo):

@Kainui

OpenStudy (kainui):

Nope, you have to square it first then multiply after, not the other way around. That's just the rules.

OpenStudy (jkbo):

well thats 80

OpenStudy (kainui):

-4 * 25 is not 80, try again.

OpenStudy (jkbo):

-100

OpenStudy (kainui):

got it, now just add them all up and you should be good.

OpenStudy (jkbo):

so -100 + (-15) + 20

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

\(y = -4x^2-3x +4\) Only plug in a 5 where there is an x.

OpenStudy (anonymous):

\(-4x^2 = -4(5^2) = -100\) That's correct. \(-3x = -3(5) = -15\) That's correct. What's left?

OpenStudy (jkbo):

4 is left @gypsy1274

OpenStudy (anonymous):

OK. so add them all together.

OpenStudy (jkbo):

I got -111

OpenStudy (anonymous):

Yes.

OpenStudy (jkbo):

thank you all

OpenStudy (anonymous):

Your welcome.